### v1

# Towards a Robust Transport Network With Self-adaptive Network Digital Twin

Abstract
[…]
keywords:
Concept drift, graph neural network (GNN), lifecycle management (LCM), twinning.
journal: Computer Networks
[…]
[lasse]organization=LASSE, Federal University of Pará,addressline=Belém, postcode=66075-110, country=Brazil
[…]
[ufg]organization=Institute of Informatics, Federal University of Goiás,addressline=Goiânia, postcode=74001-970, country=Brazil
[…]
[unisinos]organization=Applied Computing Graduate Program, University of Vale do Rio dos Sinos,addressline=São Leopoldo, postcode=93022-750, country=Brazil
[…]
[ericsson]organization=Ericsson Research,state=São Paulo, addressline=Indaiatuba, postcode=13337-300, country=Brazil
1 Introduction
[…]
NDT is a platform designed to help the management and automation of physical network communication tasks based on interactions with its virtual counterpart, also called virtual twin (VTwin), at a certain periodicity and fidelity [1, 2]. This relation allows the network operator to emulate different scenarios in network planning to perform a what-if analysis [3, 4]. This analysis aims to evaluate the impact of management changes in the network infrastructure when these tests are impossible due to logistic issues or costs that make this impossible, such as in production environments [5, 2]. For that reason, this platform is considered an enabling technology to optimize 5th and 6th generation (5G/6G) network communication [2], to provide improved solutions to problems such as slicing management [6, 7], routing optimization [8, 9], resource power allocation [3], and admission control [10].
Using the NDT platform in a production environment can create several challenges, particularly concerning the long-term accuracy of the digital twin (DT). It is essential to account for variability in incoming data, ensuring that the VTwin continually adapts to reflect real-world changes. When we consider a VTwin neural network model to achieve this reliability, such as a graph neural network (GNN[11], it is fundamental to have a well-defined lifecycle management (LCM) that organizes all phases, such as the planning, training, and operational of the NDT platform [12, 13, 14]. When a deployed NDT is used during the operational phase, an important research question arises concerning synchronization between the VTwin and the physical twin (PTwin). Synchronization issues are widely explored in the general context of DTs, considering applications at factory workstations, industrial Internet of Things, power systems, etc. [15, 16]. However, these questions are rarely explored considering the specific scenarios of a transport network NDT based on GNN. This context presents opportunities to explore particular challenges related to synchronization between twins, such as methods to achieve this and their synchronization patterns [15].
[…]
Our main contributions to this article are:
  • 1.
    A novel NDT implementation based on a GNN architecture, incorporating LCM principles during the operational phase, such as the re-deployment of an updated VTwin model when necessary.
  • 2.
    As part of our architecture, a concept drift detection module is integrated to monitor incoming network traffic features. This NDT module enables to analysis of the traffic behavior to make informed decisions about when to update the VTwin model.
  • 3.
    Relevant NDT experiments that account for environments with varying traffic conditions, including changes in traffic patterns and levels of congestion.
  • 4.
    A use case application for network management related to SLA violation prediction comparing the performance of the proposed application with and without NDT synchronization.
The remainder of this article is organized as follows: in Section 2 we compare our work with different proposed data-driven NDT in the literature, considering various levels of implementation in the transport network domain and their characteristics related to NDT supported modules, such as the synchronization one. In Section 3, we present the proposed enhanced NDT architecture, taking into account the main functional requirements of NDT, showing the essential building blocks of each, such as the VTwin architecture, the synchronization methods based on concept drift detection techniques, and complementary modules. In Section 4, we detail the experiments conducted to evaluate the proposed methods for twin synchronization and present the key evaluation results across various topologies and traffic patterns, and a use case with SLA monitoring. Finally, Section 5 summarizes the main contributions of this work and outlines directions for future research.
2 Related Work
Table 1: Comparison between NDT characteristics of the related works and our proposed architecture with support to concept drift.
Authors Closed-loop
Concept drift
detection support
Twin
synchronization
NDT LCM
elements
Rusek et al., 2020 [8]
Happ, et al., 2021 [19]
Wang et al., 2022 [6]
Güemes-Palau et al., 2022 [20]
Ferriol-Galmés et al., 2022 [21]
Saravanan et al., 2022 [22]
Lai et al., 2023 [23]
Farreras et al., 2023 [24]
Messaoudi et al., 2023 [10]
Shin et al., 2023 [25]
Yu et al., 2023 [26]
Ferriol-Galmés et al., 2023 [11]
Zalat et al., 2024 [27]
Modesto et al., 2024 [28]
Farreras et al., 2025 [7]
Aben-Athar et al., 2025 [4]
Our work
Recent studies point DTs to an important technology for the 5G/6G networks [29, 30, 31, 32, 33]. This aligns with their need to establish reliable environments for developing and testing their critical parts, such as wired environments [10], where DTs can provide excellent support. However, during this wave of adoption, the intersection of DT and networking—still an emerging area of study—led to a diverse range of approaches. Many of these did not align fully with the definitions of an NDT as systematized by standards organizations [34, 35] and industry stakeholders [2, 36]. Consequently, related work found in the literature presents limited analyses of practical aspects. For example, regarding the system lifecycle or the twinning rates, the latter refers to the rate of synchronization between the PTwin current state and its corresponding representation in the VTwin [37].
More specifically, much of the literature regarding NDTs of the transport domain focused on using machine learning (ML)/surrogate model-based methods to generate the VTwin. The increase in popularity of this DT approach, which can also be called data-driven [12], can be traced back to the work of Rusek et al. [8], which established the usage of a GNN model called RouteNet for predicting path-level metrics in routing optimization and network upgrade use cases. The usage of this model was further investigated by several works, which were built on the previously established RouteNet [19, 22, 24, 38, 11, 10, 28].
Recently, the work of Ferriol-Galmés et al. [11] utilized a modified version of RouteNet, called RouteNet-Fermi, to model network communication with various topologies, traffic distributions, and queuing configurations, enabling the prediction of packet loss, delay, and jitter. In this work, since all data is provided from a packet-level simulator, there are no two-way interactions between PTwin and VTwin. Therefore, the case presented in the article does not constitute a closed-loop NDT. Moreover, the investigations did not address the possible performance degradation of the VTwin in cases of data and/or concept drifts [39]. In this context, the cases presented in this work and several others built on RouteNet [19, 22, 24, 38, 11] did not address the robustness aspect of an NDT, as they did not explore its lifecycle either.
The exception to this lack of closed-loop implementation in works based on RouteNet is the works in Messaoudi et al. [10], Zalat et al. [27], and Aben-Athar et al. [4], which are, to the best of our knowledge, the first closed-loop environments considering physical layers interacting with an artificial intelligence (AI) model. The works in [10, 4] feed an application module tailored to operate over the physical counterpart, defined by virtual switches on Mininet [40]. The work in  [27] took the same approach, but considered a small-scale physical testbed. However, a noteworthy aspect of this work is that there are elements for twin synchronization, which are based on a trigger-based online learning approach. For the other two works, despite closed-loop interactions between the physical and virtual domains, there is no DT lifecycle management controlling the rate at which both parts synchronize and update their respective states, for instance, through a retraining process. In this context, the authors in [10, 4] ran the traffic only once, and after completing it, they performed no further operations to detect whether the VTwin performance would remain acceptable.
In addition to these, other works in the literature either proposed purely GNN-based approaches, such as FlowDT [21] or GNNetSlice [7], or adopted different ML models as their VTwins. For example, Lai et al. [23] adopted a model named eConvLSTM, Shin et al. [25] used GNN along with Long Short-Term Memory (LSTM) and a data extraction approach, [26] which uses dynamic graph attention network (DGAT). However, from these other works, only Shin et al. [25] and Yu et al. [26] presented investigations on the NDT LCM, while none presented concept drift implementations.
Finally, in terms of the datasets used in these works, there is a noticeable concentration of studies [6, 22, 7, 25, 24] based on specific datasets derived from the OMNeT++ simulator [41]. This trend presents advantages and limitations. On the one hand, using datasets from a common source provides a stable baseline, enabling fair comparisons across different approaches in terms of generalization performance. On the other hand, relying solely on pre-built datasets limits the exploration of more realistic aspects of NDT, such as the mentioned closed-loop operations or digital twin synchronization, as there is no interaction with a real or emulated environment.
Given the current state of the literature on twin synchronization in the transport domain, Table 1 summarizes the works discussed in this section based on four primary characteristics: (i) closed-loop operation, referring to the interaction between the VTwin and PTwin; (ii) support for concept drift detection to handle traffic variability; (iii) the presence of twin synchronization mechanisms, regardless of whether they explicitly use concept drift detection; and (iv) consideration of NDT LCM elements, such as the planning, training, or operational phases. The comparative analysis in Table 1 emphasizes complementary and essential features of an NDT considered in this work, rather than focusing solely on improving the VTwin model itself, as is common in prior works. In this context, our objective is not to compare different VTwin models, but rather to consider a given VTwin model and examine how our proposed architecture aims to preserve its predictive performance through twin synchronization strategies
3 Enhanced Network Digital Twin
The proposed enhanced architecture is divided into the modules illustrated in Figure 1 and shares some inheritance from standardization bodies [34, 35] and industry players [2]. Following a bottom-up description, the first module corresponds to the PTwin defined by the switches in the transport network. Above the PTwin, we have the software-defined networking (SDN) controller and its related interfaces, north and southbound. Next to it, we have the VTwin, which in our solution is implemented using a GNN model. It provides a digital representation of the relationship between network traffic and topology links. This model plays a crucial role in estimating the total per-flow delay in the PTwin. The following module is the data management, which stores useful traffic database features (see Table 2). This module maintains a unified repository of transport network data, which is essential for feeding the VTwin with relevant features in the retraining process. Additionally, this module includes the necessary functions to collect and store traffic data, ensuring seamless integration with the overall system.
Figure 1: Enhanced NDT architecture to construct a reactive approach against traffic variability. This self-adaptivity is achieved through the proposed implementation of the NDT synchronization functions (in blue), including the concept drift detector and VTwin model management.
Operating in the background to enable communication between these NDT modules, we have the interfaces defined by the SDN controller, which allow indirect data exchange between the VTwin and the controller, intended to apply any necessary modifications in the PTwin, and indirect data exchange between the database collector and the PTwin to fulfill this database with updated information from the physical environment. With this closed loop of control and traffic data between both twins, the VTwin predictions can be used in high-level applications, either with the involvement of a network operator, as defined by Aben-Athar et al. [4], or autonomously, without human supervision, as demonstrated by Messaoudi et al. [10]. In this article, we assume an SLA monitoring application. Complementary to this closed loop, the main contribution relies on the modules related to the synchronization of the NDT modules, which are responsible for synchronizing the status of the transport network of both domains in a reactive way. Along with the concept drift method, this module monitors if the incoming traffic that reaches the VTwin remains similar. Once changed and detected by the drift detector, this module has the capability to trigger a retraining process, since this change can also mean that the VTwin model is no longer predicting the total per-flow delay with an acceptable error performance. Finally, the source code of the proposed closed-loop architecture, its integration with concept drift techniques, the datasets used, and the evaluation methodology will be available in a public Git repository. [^1] https://github.com/lasseufpa/robust-ndt
3.1 Physical twin
As the PTwin, we considered a transport network topology modeled as a 2-tuple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. In this graph, $\mathcal{V}$ denotes the set of nodes, each representing a network switch, while $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ defines the set of edges (i.e., links) that connect pairs of nodes. Within this network, a traffic generator can produce a set of flow traffic instances denoted as =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{t},\dots,f_{T}\}$, with a total of $T$ flows. Each flow instance consumes network resources, such as link bandwidth, by applying variable load intensities along a specific path composed of links =$\mathcal{E}^{\prime}=\{e_{1},e_{2},\dots,e_{j},\dots,e_{J}\}$, where $J$ denotes the flow length. In this case, we define each flow instance as a tuple $f_{t}=(\mathbf{x}_{f_{t}},O,D,\mathbf{j})$, where $\mathbf{x}_{f_{t}}\in\mathbb{R}^{5}$ is a feature vector that characterizes the flow’s behavior. In this sense, each feature vector $\mathbf{x}_{f_{t}}$ is associated, by a mapping rule $h$, to a real number $y$ defined as the total per-flow delay, such that $h:\mathbf{x}_{f_{t}}\rightarrow y$. Furthermore, $O$ and $D$ represent the source and destination nodes ($\{O,D\}\in\mathbb{Z}^{+}$); and $\mathbf{j}$ is a vector of link indices that the $t$-th flow traverses. Similarly, each link instance is defined as a tuple $e_{j}=(\mathbf{x}_{e_{j}},\mathbf{i})$, where $\mathbf{x}_{e_{j}}\in\mathbb{R}^{2}$ is a feature vector describing the link’s characteristics, and $\mathbf{i}$ is a vector of flow indices representing the flows traversing that $j$-th link. The complete list of features for both flow and link instances is provided in Table 2. Furthermore, the average traffic rate $x_{t}\in\mathbf{x}_{f_{t}}$ at time $t$ is also described by a probability distribution $P_{t}$, such that $x\sim P_{t}$.
Table 2: Flow and link features that are extracted to be used as input feature for training and testing the VTwin model.
Feature Unit Type
Average traffic rate $\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$ Flow
Path propagation delay $\mathrm{s}$ Flow
Flow length Flow
Average packet sent Flow
Average packet loss $\mathrm{p}\mathrm{a}\mathrm{c}\mathrm{k}\mathrm{e}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$ Flow
Capacity $\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$ Link
Link load % Link
3.2 Database management and SDN controller
The database used for the decision-making process is divided into three storage types. The first one, called the traffic database, is related to the data stream that was filtered after the traffic generator created it and will serve as input for the VTwin model to perform the necessary delay prediction, i.e., the total per-flow delay label is unavailable. The second type of database is related to the examples that should be used to retrain our VTwin model, and it is called labeled database, as we manipulated the ground truth delay to update the VTwin model weights when necessary. The last storage refers to the weights database, aimed to store the historical VTwin model weights during the NDT operation.
To store these raw data, we are also considering data collection functions to discard unnecessary packet-level traffic data, such as the Internet Protocol (IP) addresses of the server and client, or transmission time timestamps. Therefore, we aim to store only quality of service (QoS) related features that we can extract from the PTwin environment and that are useful for the VTwin prediction task. It is important to note that we assume that the data required for the retraining process has already been collected, labeled, and stored in our database. Furthermore, issues related to synthetic data generation or real-time data collection for retraining are considered outside the scope of this work.
The interfaces employed in the proposed architecture (Figure 1) enable the indirect exchange of control data between both twins, with the SDN controller serving as the network operating system [42]. To allow this interaction, we utilize the SDN controller as the central bridging tool, leveraging its available application programming interface (API). Specifically, another feature stored in the traffic and labeled database consists of the vectors $\mathbf{i}$ and $\mathbf{j}$, which represent, respectively, the links traversed by a given flow and the flows that have traversed a given link. Consequently, one of the critical functions of the interface module is to work with the data collection’s path collector, which identifies and records the routes taken between each source–destination node pair. The path collector operates based on a specific node identifier, which is a 16-character hexadecimal string derived from the PTwin. Therefore, given that flow traffic was initialized between two pairs of nodes, it is possible to consult, by the controller REpresentational State Transfer (REST) API, the path of each flow. In this case, it is assumed that the SDN controller operates in a reactive forwarding way, that is, considering a forwarding decision and route establishment on demand, working only when a packet should be sent. Moreover, the shortest path defines the routing table of each switch in the network.
3.3 Virtual twin
Considering the adopted NDT architecture, the VTwin assumes a role in mimicking key aspects of its physical counterpart. In this sense, the VTwin is based on a GNN, a type of deep learning model. Accordingly, the VTwin is responsible for abstracting the most complex organization of the transport network and synthesizing its key aspects to evaluate the interest QoS parameter, keeping a certain level of fidelity in the physical-virtual relationship. For the objective of estimating the total flow delay, it is essential to model the relation between the flow traffic applied to propagate information through the wired medium and the links used for it, as shown in the generic graph in Figure 2. This modeling step is fundamental, as the VTwin input is not the original graph representing the transport network but rather the hypergraph derived from it. The relation between the flows and links of Figure 2 creates an essential circular dependency that is explored to generate the heterogeneous graph used as input to the VTwin model. In this case, we can see that flow $f_{1}$ depends on links $e_{1}$, $e_{2}$, and $e_{5}$. Similarly, the flow $f_{2}$ has a dependency on the links $e_{3}$, $e_{4}$, and $e_{5}$. Therefore, these relationships generate the VTwin input graph depicted in Figure 3.
Figure 2: Links and flows relation in a generic network topology.
Figure 3: Heterogeneous graph, considering flow and link node types, generated from the transport network of Figure 2.
Regarding the architecture of the VTwin, we have a GNN based on modified RouteNet-Fermi with a self-attention mechanism [28], as illustrated in Figure 4. This model processes flow and link characteristics through message passing across two gated recurrent unit (GRU)-based layers. The aggregated messages are used by a readout function defined by a multilayer perceptron (MLP) to compute occupancy and estimate the total per-flow delay.
Figure 4: Flow and link features from generic topology processed by the VTwin model.
The embedding initialization has the purpose of processing the node features $\mathbf{x}_{f_{t}}$ and $\mathbf{x}_{e_{j}}$ from a heterogeneous graph obtained from the relationship between the links and the flows, that is, the $f_{t}$ flows that passed by several links and the $e_{j}$ links that received each flow [28]. Since this graph has two types of nodes with different feature vectors, specific embedding initialization functions are used to convert the features of each node type from their original dimension to high-dimensional space $\mathbf{h}_{f_{t}}\in\mathbb{R}^{m},\mathbf{h}_{e_{j}}\in\mathbb{R}^{n}$, where $m$ and $n$ are model hyperparameters, respectively, the size of the flow and link embedding vectors.
To obtain these high-dimensional vectors for each node type, two MLP layers are used as an initialization function, both with one hidden layer and having the output dimension defining the embedding size, $\|\mathbf{h}_{f_{t}}\|=m$ and $\|\mathbf{h}_{e_{j}}\|=n$. Thereby, this embedding initialization is defined by equations:
$\mathbf{h}_{f_{t}}=\text{MLP}(\mathbf{x}_{f_{t}},\theta_{f_{t}}),$ (1)
and
$\mathbf{h}_{e_{j}}=\text{MLP}(\mathbf{x}_{e_{j}},\theta_{e_{j}}),$ (2)
where $\theta_{f_{t}}$, $\theta_{e_{j}}$ are the learnable parameters of each initialization function.
In the message-passing phase, these high-dimensional vectors are processed iteratively during $k\leq K$ (which $K$ is another model hyperparameter) time steps by the message function $\psi(\cdot)$ and updated by functions $\phi_{\textrm{f}}(\cdot)$, $\phi_{\textrm{l}}(\cdot)$, as defined as follows [43]:
$\mathbf{h}_{f_{t}}^{k+1}=\phi_{\textrm{f}}\left(\mathbf{h}_{f_{t}}^{k},\psi(\mathbf{h}_{f_{t}}^{k},\mathbf{h}_{e_{j}}^{k})\right),$ (3)
and
$\mathbf{h}_{e_{j}}^{k+1}=\phi_{\textrm{l}}\left(\mathbf{h}_{e_{j}}^{k},\bigoplus_{p\in\mathcal{N}_{e}}a(\mathbf{h}_{e_{j}}^{k},\mathbf{h}_{p}^{k+1})\psi(\mathbf{h}_{e_{j}}^{k},\mathbf{h}_{p}^{k+1})\right),$ (4)
where $\bigoplus$ is an aggregation function, $a(\cdot)$ the attention score, and $\mathcal{N}_{e}$ is the neighborhood of $e$-th link node. Moreover, this model treats the message function as a GRU layer and the auxiliary assignment operation as an update function.
The output of the message-passing module computes the occupancy $\mathcal{O}$ through a readout function, implemented using an MLP with two hidden layers. The MLP produces a single-neuron output that yields the occupancy rate for each link in the network, as defined by the following equation:
$\mathcal{O}=\text{MLP}(\mathbf{h}_{f_{t}}^{k+1},\theta_{\textrm{o}}),$ (5)
where $\theta_{\textrm{o}}$ are the learnable parameters of the readout MLP layer, considering that a flow set passes through $J$ links, generating the occupancy mentioned above $\mathcal{O}$, the predicted total per-flow delay $\hat{y}$ is evaluated by:
$\hat{y}=\sum_{j=0}^{J}\left(\dfrac{\mathcal{O}_{j}}{\mathcal{E}_{j}^{\prime}}\right).$ (6)
3.4 NDT Synchronization
A well-defined LCM provides the NDT platform with the essential capabilities to function effectively in a production environment. It defines how the NDT should behave under various conditions, especially during the operational phase [14], whether those conditions are routine or exceptional. Consequently, all LCM phases can encompass different modules of the architecture presented in Figure 1, and each of them has a specific role and lifespan to ensure the correct operation of the NDT tasks [14]. Since each phase has a different lifespan, some modules shown in Figure 1 may not exist in particular phases. For instance, during the planning stage of the VTwin, the PTwin is not yet required. At this point, the VTwin can be trained using pre-existing datasets [12], considering related features from the target network environment, eliminating the need for real-time data integration. As a result, only the VTwin and selected components of the data management module are active. This phase closely resembles the process of training an AI model for predictive tasks using offline data, where the focus is on optimizing model parameters and hyperparameters based on generalization performance. Noteworthy is that the early stages of NDT for the transport network, as seen in the related works, had a specific focus in this phase.
As the NDT lifecycle progresses, the training phase begins to incorporate physical components of the SDN infrastructure, including the interfaces connected to the SDN controller. In this context, training refers to the process of evaluating the VTwin prior to its deployment in a production environment. This phase enables the identification and mitigation of potential critical issues that could impact the proper functioning of the NDT platform. Finally, in the operational phase, we employ two submodules that monitor the synchronization states of the NDT: concept drift detector and VTwin model management. The relationship in terms of input and output during the NDT operational phase is defined by Figure 5. These submodules are essential as they ensure the NDT robustness against traffic variability after NDT planning and deployment, which is intimately associated with the degree of synchronization between the VTwin and PTwin [15, 33]. This twinning rate will determine the frequency at which the NDT will start the process of synchronizing the network states of the VTwin [37].
Figure 5: Sequence of steps in the operational phase to synchronize the PTwin states in the VTwin.
Concept drift detector
In a data-driven supervised learning approach, the synchronization process is associated with a retraining operation, which can be triggered based on the input or output of the neural network model. This choice is essential, as the synchronization between the VTwin and PTwin will differ significantly depending on the target variable considered in the concept drift detector process. In the first case, synchronization is driven by the distribution pattern of each input feature, such as the flow of traffic passing through a set of links. In this sense, when decisions regarding concept drift are made solely based on the input data, the approach is referred to as a distribution-based method [44]. In contrast, when the correct labels are also considered, the method is classified as performance-based [44]. In the latter approach, in the context of this problem, synchronization depends on the label, specifically the total per-flow delay. However, using this approach, under realistic conditions, detecting the aging of a neural network model is not straightforward, as the ground truth delay is not readily available for direct comparison with VTwin’s predictions. Therefore, our synchronization approach employs a univariate distribution-based method that monitors the flow traffic rate to detect concept drift during the VTwin operation, thereby eliminating the need to rely on ground truth delay to trigger the synchronization process through VTwin retraining. In operational terms, to reach these synchronization steps, we need to pass through different NDT modules, as defined by the flowchart in Figure 5.
The process starts with user traffic originating from the transport network. This raw traffic data is captured by the relevant network functions and stored in the traffic database. The data collector then extracts a key feature for use by the concept drift detector: the average traffic rate of each flow. This feature is selected because it effectively captures the primary characteristic of user traffic, its average throughput. Thereby, using a univariate covariate-based method as a concept drift detector, we can identify changes in the input distribution of features over time when $P_{t}(x_{f_{t}})\neq P_{t+w}(x_{f_{t+w}})$, where $t$ represents a given time instant, $\{x_{f_{t}}\in\mathbf{x}_{f_{t}},x_{f_{t+w}}\in\mathbf{x}_{f_{t+w}}\}$ is a feature in time instant $t$ and $t+w$, in this case, the average traffic rate, and $w$ are the sizes of the time window. In this context, our concept drift detector serves as a monitoring agent that detects changes in traffic patterns as incoming traffic is processed. This process allows us to indirectly estimate when the total per-flow delay distribution will begin to shift, signaling the expected degradation in VTwin performance and triggering retraining.
VTwin model management
When concept drift is detected, the concept drift detector notifies the VTwin model management submodule with a warning to start retraining using data already available in the labeled database. Upon the start of the retraining process, the VTwin model management locks the trigger to prevent the drift detector from issuing duplicate warnings for the same drift event previously addressed until the training is finished. Upon completion of the retraining process, the VTwin model management sends a control message to the VTwin, archives the previously trained model weights in a dedicated database, and updates the VTwin model with the newly learned weights.
3.5 Network application
The network application is organized to use the predictions provided to the VTwin to perform adjustments in the PTwin considering applications that possibly will change the state of the PTwin, such as routing optimization or flow admission control or related to the management of the network, such as SLA monitoring for what-if analysis [6] or emergency preparedness [45]. To achieve this, various submodules can be employed within these applications, including, but not limited to, a recommendation system to propose different configurations in the PTwin, a policy evaluation module to evaluate the configurations suggested by the recommendation system, and a network management submodule designed to provide application-level metrics from the PTwin.
Recommendation system
In scenarios where the network application aims to optimize certain states of the PTwin, this module incorporates auxiliary tools, such as a recommendation system, that assist the network operator by leveraging outputs from the VTwin. This recommendation system can generate new policies that can be potentially applied to the PTwin. For example, in routing optimization, the recommendation system, which can be based on a genetic algorithm [27] or a reinforcement learning model [4], may propose new routing policies based on total per-flow delay metrics provided by the VTwin, aimed at optimizing the QoS metrics on the transport network. Once suggested, this recommendation is sent to the policy evaluation submodule, to be evaluated by a network operator, if a man-in-the-loop NDT is being used [2].
Policy evaluation
In a man-in-the-loop architecture, a policy evaluation submodule is essential to prevent network modifications without human oversight, an important safeguard in production environments, particularly in architectures built on 5th generation (5G) networks [10]. This is especially critical given the interdependence of network components, such as the reliance of the radio access network on the transport network. The role of this submodule is to consolidate all relevant information and assess the potential impacts on the network, enabling informed decision-making by the operator. Once a policy is evaluated, the final decision on its implementation rests with the operator. If approved, the corresponding control data is transmitted to the SDN controller via the northbound API and subsequently applied to the PTwin, thus completing the NDT loop.
Network management
When the application serves as a network management tool for what-if analysis, the task becomes conceptually simpler, as it does not require direct interaction with the PTwin to implement changes. Instead, its main objective is to provide the network operator with insight into the possible outcomes of specific scenarios or conditions [45], which can be useful in network planning tasks. This work explores a particular use case in the Experimental Results section, focusing exclusively on the network management submodule to provide SLA violations in the network [46, 47]. The selected application assesses compliance with flow SLAs, specifically detecting violations related to the packet delay budget (PDB) assigned to each packet, for example, if the total per-flow delay exceeds the PDB assigned for this flow. So, for a given set of flows that traverse the transport network, this application can estimate the number of SLA violations, relying on the total per-flow delay predictions provided by the VTwin.
4 Experiments
The PTwin framework integrates both emulation and simulation components to evaluate the proposed architecture. Using Mininet with open network operating system (ONOS) controller, the emulated component focuses on topology organization, while the simulated component handles traffic generation. For emulated network topologies, we use three transport networks: the PASSION project topology [48], the 5G-Crosshaul topology [48], and the Germany-50 topology [49], comprising networks with 128, 51, and 50 nodes, respectively, and each with 224, 88, and 63 links. Moreover, we employ the Distributed Internet Traffic Generator (D-ITG) using a client-server model over the Transmission Control Protocol (TCP) and User Datagram Protocol (UDP) protocols to generate realistic traffic patterns. To do so, we used a computer equipped with Intel® Core i7-10700F CPU @ 2.90 GHz, 128 GB of RAM, and a storage capacity of 9 TB. This setup enables the incorporation of various traffic distributions within an emulation environment that is orchestrated using a suite of tools, summarized in Table 3, to ensure accurate and reproducible testing conditions.
Table 3: Setup configuration used to build our PTwin.
Tools Version
Operating system Ubuntu 20.04
ONOS 2.6.0
Mininet 2.3.0
Open vSwitch 2.13.8
OpenFlow 1.0
D-ITG 2.8.1
We consider different distributions of packet size (in bytes) and traffic rate (in $\mathrm{M}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$) distributions, without loss of generality, used to model internet traffic [50, 51], such as Poisson, Exponential, Uniform, and Deterministic, with the presence of congestion. The parameters for each are defined in Table 4, considering the minimum and maximum values $\mathcal{U}(512,1024)$ for the uniform distribution, the mean for the exponential $\mathcal{E}(1024)$ and $\mathcal{P}(2048)$ for a Poisson distribution, and a deterministic pattern with traffic congestion defined by a constant value $K=512$ for a constant packet size in traffic generation. For the packet rate, we considered the same distribution as defined for the packet size.
Table 4: Packet size distributions used to simulate user traffic.
Traffic pattern
Protocol
Packet size
Uniform TCP $\mathcal{U}(512,1024)$
Exponential TCP $\mathcal{E}(1024)$
Poisson TCP $\mathcal{P}(2048)$
Deterministic UDP $K=512$
The datasets created from this traffic generation for a possible retraining process are presented in Table 5, characterized by the four distribution patterns of packet size and traffic rate, the type of topology, and the number of training examples. The training examples in this case refer to the number of samples that are to be used as input-output in the model retraining.
Table 5: Type of labeled datasets stored used to retraining the VTwin.
Traffic pattern
Topology
Training
examples
1 Exponential 5G-Crosshaul 121 400121\,400121 400
Germany 129 600129\,600129 600
PASSION 129 600129\,600129 600
2 Poisson 5G-Crosshaul 234 501234\,501234 501
Germany 257 401257\,401257 401
PASSION 243 001243\,001243 001
3 Uniform 5G-Crosshaul 347 002347\,002347 002
Germany 378 802378\,802378 802
PASSION 355 002355\,002355 002
4
Deterministic
5G-Crosshaul 447 803447\,803447 803
Germany 440 003440\,003440 003
PASSION 478 203478\,203478 203
Considering these emulated environments into account, we propose an experiment using topologies that experience a sudden concept drift in traffic characteristics. We introduced an artificial drift at specific time intervals by altering the packet size distribution. As a result, the traffic follows the sequence of packet size and traffic rate patterns described below:
Start $\rightarrow \mathcal{E}$rightarrow $\mathcal{P}$rightarrow $\mathcal{U}$rightarrow $K$rightarrow End
For each topology, we analyzed different time intervals of network traffic. Specifically, the total times in seconds for 5G-Crosshaul, Germany, and PASSION are 25 10325\,10325 103 $\mathrm{s}$, 24 60324\,60324 603 $\mathrm{s}$, and 26 70326\,70326 703 $\mathrm{s}$, respectively. Moreover, the experiment was carried out considering that we initially trained the VTwin model using traffic flows characterized by an exponential packet size distribution. The first abrupt drift is introduced by switching the packet size distribution to a Poisson model. This changing distribution behavior persists until the final phase, where the traffic pattern transitions to a constant packet size distribution, accompanied by the presence of traffic congestion.
Finally, we considered as a metric the NMSE to evaluate performance throughout these concept drifts, defined as:
$\text{NMSE}=10\log_{10}\left\{\dfrac{\sum\limits_{t=1}^{T}\left(y_{t}-\hat{y}_{t}\right)^{2}}{\sum\limits_{t=1}^{T}y_{t}^{2}}\right\},$ (7)
where $\hat{y}_{t}$ and $y_{t}$ represent the $t$-th estimated total per-flow delay and the corresponding ground truth total per-flow delay, respectively, across a total of $T$ examples in the test dataset. To compute the average, we used a window size of 100 consecutive examples over time. While other window sizes could be adopted, using 100 samples strikes a balance: it yields a sufficiently narrow confidence interval for the mean and enhances sensitivity in terms of NMSE to short-term variations, particularly important in the presence of abrupt concept drift.
Using this environment, we present an experiment that considers a distribution-based detector aimed at detecting input changes. We configured the concept drift detector using the Kolmogorov-Smirnov windowing (KSWIN) method [52], as implemented by the river library [53]. To work correctly, this method requires three hyperparameters: the significance level, which is a sensitive parameter that controls the probability of false positives, represented by a value of 0.001, a window size, and a buffer for the most recent samples. The sizes of these components vary across different topologies and were determined empirically.
For the VTwin training, we used TensorFlow 2.14.0 and maintained consistency in all training parameters in all experiments. Specifically, we used a learning rate of 0.0010.0010.001, trained for 50 epochs, applied Z-score normalization for feature scaling, adopted the Adam optimizer, and used Mean Absolute Percentage Error (MAPE) as the loss function. Moreover, as the model hyperparameters, we used: $T=12$ and $||\mathbf{h}_{f}||=||\mathbf{h}_{l}||=16$. This uniform configuration was chosen to ensure that any observed differences in model performance could be attributed solely to variations in the input data.
Experimental Results
We have the experiments from the KSWIN results to evaluate the proposed synchronization elements, considering the three emulated topologies and the average traffic rate in $\mathrm{M}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$ as the feature to be tracked. Figure 6 illustrates the operation of the KSWIN detector and the location of each detected drift. For each network topology, we selected a specific window size (in seconds): 6800 for 5G-Crosshaul, 7000 for Germany, and 6800 for the PASSION topology. Although alternative window sizes could be explored, determining the optimal size to minimize false positives falls outside the scope of this work.
Figure 6: Drift detection pattern considering 5G-Crosshaul, Germany and PASSION topologies and four type of traffic patterns.
In the evaluated environments, smaller window sizes generally result in a higher likelihood of false positives, as illustrated in Figure 7, which analyzes various window sizes in all topologies. The curves in this figure exhibit a consistent pattern and converge in the number of detected concept drifts as the window size increases. This behavior can be attributed to the timing of the concept drift events. For example, the first concept drift, shown in Figure 6, occurs approximately in the second 7000, across all topologies, coincident with the point at which the curves in Figure 7 end converging.
Figure 7: Relationship between the window size and the number of concept drift detection on the three topologies.
In this experiment, we can observe that concept drift was effectively detected in all existing deviations due to changes in the traffic pattern. However, there were different delays, considering the time when the concept drift occurred and when the detection was made. Considering the effects of these delayed drift detections, we can see in Figure 8 the VTwin performance over time considering the proposed synchronization elements in comparison with the same VTwin without them in terms of NMSE.
Figure 8: Validating proposed architecture on 5G-Crosshaul topology. The blue line represents the VTwin performance with the proposed synchronization elements, while the gray dashed line represents the performance without it.
Before the first concept drift is detected, the VTwin operates in a typical manner, as seen in the related work: a GNN model already trained to generalize to unseen datasets and, based on its predictions, performs the required analysis within the proposed application. However, after the first concept drift occurs, the scenario changes due to traffic variability, leading to model degradation, evidenced by some sample windows exhibiting NMSE values near 0 $\mathrm{d}\mathrm{B}$. Despite this decline, the VTwin model remains stable, even without retraining, as the underlying data distributions, such as Poisson and Exponential, are still related. This behavior changes after the second concept drift, as the input data distribution changes to a packet size with a distinct pattern, requiring retraining, which is evident in the accentuated NMSE difference between the model with and without retraining. In the last concept drift, the VTwin becomes obsolete without a retraining process as the input data distribution shifts to a UDP connection characterized by constant packet size and the absence of congestion control. At this point, retraining becomes extremely necessary, as demonstrated by the pronounced performance gap between the models with and without retraining.
The results for the other two topologies aim to validate the purpose synchronization method considering topologies with different characteristics, for instance, in the PASSION topology that has 128 nodes and three times more links than the 5G-crosshaul. In this way, Figure 9 also shows consistent results in both topologies as those observed in the 5G-Crosshaul topology, specifically in the last drift detected by the concept drift detector, highlighting the same observations related to mandatory retraining due to poor VTwin prediction performance.
Figure 9: Evaluation performance of proposed architecture considering Germany and PASSION topologies.
[…]
Table 6: NMSE performance in $\mathrm{d}\mathrm{B}$ before and after concept drift, considering the NDT operation with and without the proposed synchronization module implementation
Topology Drift ID
Pre-drift avg.
NMSE w/o sync.
Post-drift avg.
NMSE w/o sync.
Pre-drift avg.
NMSE w/ sync.
Post-drift avg.
NMSE w/ sync.
5G-Crosshaul 1 20.72-20.72- 20.72 10.84-10.84- 10.84 20.72-20.72- 20.72 $\mathbf{-14.47}$
2 10.84-10.84- 10.84 5.42-5.42- 5.42 14.47-14.47- 14.47 $\mathbf{-11.22}$
3 5.42-5.42- 5.42 7.787.787.78 11.22-11.22- 11.22 $\mathbf{1.43}$
Germany 1 19.98-19.98- 19.98 12.11-12.11- 12.11 19.98-19.98- 19.98 $\mathbf{-20.10}$
2 12.11-12.11- 12.11 1.50-1.50- 1.50 20.10-20.10- 20.10 $\mathbf{-14.85}$
3 1.50-1.50- 1.50 25.1025.1025.10 14.85-14.85- 14.85 $\mathbf{-5.73}$
PASSION 1 18.44-18.44- 18.44 6.62-6.62- 6.62 18.44-18.44- 18.44 $\mathbf{-19.17}$
2 6.62-6.62- 6.62 0.470.470.47 19.17-19.17- 19.17 $\mathbf{-8.04}$
3 0.470.470.47 38.7038.7038.70 8.04-8.04- 8.04 $\mathbf{5.83}$
Use case with SLA violation monitoring
The use case scenario used to evaluate the proposed NDT synchronization, considering a high-level metric related to SLA monitoring tasks, was carried out using only data coming from 5G-crosshaul topology. In this sense, we evaluated the same scenario presented in Figure 8, but considering high-level metrics such as the accuracy in the SLA violation predictions. Thus, we assume that each flow has a specific PDB value that was assigned empirically, based on the characteristics of the 5G-crosshaul topology, such as the propagation delay of each link.
Regarding the experiments, we consider a total of 25 10025\,10025 100 flows traversing the network, each of which must be classified by the SLA monitoring application as compliant or in violation with the PDB, based on the delay predicted by the VTwin. The SLA violation in this case is characterized when a per-flow delay exceeds the PDB for a given path. From the total of flows considered in this experiment, 536536536 flows are violating with their SLA (ground truth). Hence, the accuracy in the prediction of SLA violations is shown in Figure 10, which considers the VTwin with the NDT synchronization module and without it, and each window index corresponds to the classification of 100 flows, for a better visualization of the results in this figure.
Figure 10: SLA violations in scenario without (bottom) and with twin synchronization (upper). The orange curve show the number of actual SLA validation, and the blue curve the number of predicted violations.
As shown in Figure 10, the application exhibits the expected degradation in classification performance when the NDT synchronization module is disabled, due to the VTwin states not being updated with PTwin states. Quantitatively, this results in a misclassification of 7 7447\,7447 744 flows that are not in compliance with their respective SLAs. In contrast, when the NDT synchronization module is enabled, the number of incorrectly classified flows is significantly reduced to 3 $3$\,$3$ 333. It is important to note that part of these SLA violations also occurs during periods when the VTwin is undergoing retraining. Overall, this experiment highlights that the practical usability of an NDT strongly depends on the presence of an effective synchronization mechanism between the twins, operating at an acceptable rate, which is essential for transitioning the solution from a prototype to a production-ready platform or framework.
5 Conclusions and Future Works
[…]
The results of the proposed experiments offer valuable insights into the challenges of twin synchronization, positioning this article as a foundational step toward future developments. A key direction for future work involves progressively incorporating more realistic elements into the experimental environments. One major limitation is the gap between the current setup and real-world production environments. Although the proposed module presents a promising approach to addressing specific challenges faced by NDT systems, it does not fully resolve all the issues required for the complete feasibility of the NDT platform. For instance, while the data management module includes essential mechanisms for collecting information from the PTwin, the experiments rely on traffic generated by a traffic generator, with metrics directly obtained from simulator logs. In contrast, data collection in a production environment is significantly more complex and may require integration with auxiliary tools, such as those defined by protocols such as the Simple Network Management Protocol (SNMP) used to collect network metrics. Furthermore, the feasibility of the proposed synchronization method also depends on the ability to generate synthetic data to retrain the VTwin model when necessary, which is not addressed in this article. In this context, the use of generative AI emerges as a promising solution and represents a potential avenue for future research.
In addition to the data management tasks associated with an NDT, future research will also focus on improving the synchronization method itself, which can be pursued along two main directions. In the short term, the goal is to evaluate synchronization performance in variable transport network environments using a more robust multivariate concept drift detection approach, one that simultaneously considers multiple input features such as traffic rate, per-link load, and link capacity. In the long term, we aim to develop a more proactive synchronization strategy by integrating online learning techniques. This would reduce the latency in adapting the VTwin model parameters without waiting for an explicit concept drift detection, thereby minimizing the duration which outdated model weights are in effect.
CRediT authorship contribution statement
Cláudio Modesto: Writing — original draft, Writing — review and editing, Visualization, Validation, Software, Methodology, Investigation, Data curation, Conceptualization. João Borges: Writing — original draft, Conceptualization. Cleverson Nahum: Software, Supervision, Writing – review and editing. Lucas Matni: Software, Formal analysis. Cristiano Bonato Both: Writing – review and editing, Supervision, Visualization. Kleber Cardoso: Writing – review and editing, Methodology. Glauco Gonçalves: Writing – review and editing, Formal analysis. Ilan Correa: Writing – review and editing. Silvia Lins: Writing – review and editing, Project administration. Andrey Silva: Writing – review and editing, Project administration. Aldebaro Klautau: Writing – review and editing, Project administration, Funding acquisition.
Acknowledgments
This study was financed in part by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) – Finance Code 001, the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq); the Innovation Center, Ericsson Telecomunicações Ltda., Brazil; and Project Smart 5G Core and MUltiRAn Integration (SAMURAI) (MC-TIC/CGI.br/FAPESP under grant 2020/05127-2).
Data Availability
Data will be made available on request.